# SketchSplat: 3D Edge Reconstruction via Differentiable Multi-view Sketch Splatting

### v1

@@ 5 Experiments
Figure 3: Qualitative comparison on the ABC-NEF dataset [53]. Our method achieves the most complete results compared to the baselines.
@@
R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$ R5$\uparrow$ P5$\uparrow$
@@
5.1 Experimental Setup
Datasets.
We evaluate our method on two datasets: ABC-NEF [53, 18] and DTU [1]. The ABC-NEF dataset [53, 18] comprises 115 CAD models, each containing 50 RGB images of the target object. Each model is accompanied with a parametric edge file providing ground truth 3D edges. Following EMAP [19] and EdgeGS [7], we use 82 CAD models from ABC-NEF, excluding the ones with cylinders and spheres since smooth convex surfaces will produce silhouette edges that are not consistent in multi-views. For 2D edge image generation, we use PiDiNet [43], DexiNed [33], and our proposed 2DGS-SN as 2D edge detector. The DTU dataset [1] contains multi-view images of real-world scenes captured from forward-facing views. We follow EdgeGS [7] and evaluate on 6 DTU scenes. We use edge images generated by PiDiNet [43] and the pseudo ground truth edges provided by EMAP [19]. To generate pseudo ground truth edges, EMAP leverages the ground truth 3D surface points and filters out the points that do not project onto 2D edges.
Baselines.
We compare our method with five baseline methods, including the state-of-the-art 3D line mapping method (LIMAP [23]), and several methods based on volumetric differentiable rendering (NEF [53], EMAP [19], NEAT [50], EdgeGS [7]). Following EMAP and EdgeGS, we do not evaluate NEAT on the ABC-NEF dataset since the authors of EMAP found that NEAT often fails to train for texture-less objects.
Metrics.
Following the metrics used in baseline methods [19, 7], we evaluate Accuracy (A) and Completeness (C) in millimeters, and Recall (R), Precision (P), and F-score (F) as percentages at three thresholds (5mm, 10mm, 20mm). All of the scenes are scaled to fit within a $1\text{m}^{3}$ bounding box. See supplementary material for more details.
5.2 Results
Figure 4: Ablation study on the proposed components of SketchSplat.
Evaluation on ABC-NEF [53].
The qualitative and quantitative evaluation results on ABC-NEF dataset are shown in Fig. 3 and Tab. 1, respectively. Overall, the proposed method achieves a new state-of-the-art across most of the metrics. Notably, it is the first time to achieve both accuracy (A) and completeness (C) lower than 7.0mm, while also surpassing 90% in Recall (R5), precision(P5), and F-score (F5).
Tab. 1 shows LIMAP [23] achieves the highest accuracy and precision (P5) due to its accurate edge detection and robust line reconstruction method. However, its completeness and recall are low, as LIMAP struggles with curve reconstruction, and inconsistent edge detection leads to missing lines, as shown in Fig. 3. Our method (with 2DGS-SN) achieves comparable accuracy but much higher completeness. NEF produces visually complete results in Fig. 3(c), but performs poorly on metrics due to its fitting process relying solely on extracted 3D edge points. EMAP outperforms NEF with its unsigned distance field modeling and point shifting mechanism. However, Fig. 3(d) shows that EMAP still generates duplicated edges, misses some edges, and struggles with accurate curve fitting. EdgeGS surpasses EMAP in recall and precision at 10mm and 20mm thresholds, but the trend reverses for measurements under 5mm. This is due to the thickness of 2D edges introduces bias in smaller distance evaluations. As shown in Fig. 3, our SketchSplat achieves the highest completeness and compactness (edges are well connected) compared to all the baselines.
As shown in Tab. 1, the proposed 2DGS-SN helps boost performance of both EdgeGS and SketchSplat to a large margin, which shows 2DGS-SN overcomes the bias introduced by 2D edge detectors PiDiNet and DexiNed. It should be noted that when evaluated with all three edge detectors, our method consistently achieves higher Recall, Accuracy, and F-score than EdgeGS [7], which validates the effectiveness of SketchSplat in better aligning 3D edges with 2D edge images.
Method E-num Time (h)
NEF [53] 22.9 1:26
EMAP [19] 45.8 2:30
EdgeGS [7] 140.8 0:05
ours 44.3 0:10
Table 3: Comparison of the number of reconstructed edges (E-num) and runtime on ABC-NEF.
Evaluation on DTU.
Tab. 2 shows the evaluation results on six scenes in the DTU dataset [1]. We reproduce EdgeGS [7] using its released code and configuration (denoted as EdgeGS*) to initialize sketches for our method. As shown in Tab. 2, SketchSplat without topological operations achieves comparable results to EMAP [19] and slightly outperforms EdgeGS, demonstrating its effectiveness in correcting misalignment between 3D edges and 2D edge images. However, incorporating topological operations slightly lowers recall. This is because the merging operations aim to merge repeated edges and make the results compact, while the pseudo ground truth edges favor duplicated edges in precision (P5) and recall (R5).
Speed Analysis.
Compactness.
EdgeGS achieves higher recall than NEF and EMAP due to its strict edge fitting process, which produces numerous short edges, covering more edge regions as the number of required edges increases. Instead, our method incorporates adaptive topology control to reduce redundant edges, as shown in Tab. 3 and Fig. 4(a,b,d). This mechanism helps SketchSplat achieve higher compactness, while maintaining the highest accuracy and completeness.
5.3 Ablation
[…]
Ours 6.8 5.8 90.8 92.9 91.3 44.28
w/o 2DGS-SN 8.5 8.6 47.5 54.1 50.3 33.52
w/o Merge 6.6 5.7 91.8 92.7 91.7 140.72
w/o Filter 6.8 5.8 90.9 92.9 91.3 44.37
Table 4: Ablation on the proposed components using ABC-NEF. ‘w/o 2DGS-SN’ denotes SketchSplat trained with edge images detected by DexiNed, while the others are trained with 2DGS-SN.
Contributions.
[…]
EdgeGS 7.4 7.2 75.7 86.9 80.3 140.76
Ours (EdgeGS) 6.8 5.8 90.8 92.9 91.3 44.28
EMAP 8.8 7.9 63.5 70.4 66.3 48.71
Ours (EMAP) 6.9 6.2 88.8 90.8 89.3 31.70
Table 5: Ablation on initialization methods using ABC-NEF. ‘Ours (EdgeGS)’ and ‘Ours (EMAP)’ denote our method initialized with EdgeGS and EMAP, respectively. 2DGS-SN is chosen as 2D edge detector.
Initialization Methods.
In our main experiments, we use EdgeGS for initialization. However, we also demonstrate SketchSplat can achieve strong results when initialized with EMAP. As shown in Tab. 5, our approach consistently outperforms the initialization methods and achieves comparable results across different initialization methods.
6 Conclusion
We present SketchSplat, a method for parametric 3D edge reconstruction from calibrated multi-view images. Unlike previous approaches that rely on 3D edge point reconstruction and edge fitting, SketchSplat directly optimizes 3D edges through differentiable multi-view sketch splatting, ensuring accurate alignment with input images. A series of topological operations are proposed to adaptively refine the edge structure during training, which reduces redundancy while maintaining high accuracy. Additionally, an improved 2D edge detection method is adopted to achieve higher reconstruction quality. Experiments show that SketchSplat achieves state-of-the-art accuracy and completeness on CAD datasets while maintaining compact results and efficient training speed.
Acknowledgements
Supplementary Material of
SketchSplat: 3D Edge Reconstruction via Differentiable Multi-view Sketch Splatting
Summary.
In this supplementary material, we elaborate on the following subjects. We describe the details of the metrics utilized in the main paper in Sec. 7. More visual results (including 20 scenes in ABC-NEF dataset) are discussed and shown in Sec. 8. In Sec. 9, we provide more details of our proposed edge detector 2DGS-SN. Finally in Sec. 10, we discuss about the limitations of our method.
@@ 8 More Visual Results
9 Details of 2DGS-SN
We show the accuracy of 2DGS-SN by comparing with PiDiNet[43] and DexiNed[33]. From Fig. 5, we can see (b) 2DGS-SN produces accurate edges, where the object boundary locates at the center of the predicted edge pixels. Both PiDiNet and DexiNed detect edges with slight offsets, positioning them far from the object boundary. According to Tab. 1 in the main paper, this slight offset introduces significant multi-view inconsistency and ambiguity, resulting in degraded performance.This explains why existing methods like EdgeGS achieve much higher F10 and F20 scores compared to F5.
Figure 5: Visual comparison of performance of different edge detection methods.

### v2

@@ 5.3 Ablation
Ours 6.8 5.8 90.8 92.9 91.3 44.28
w/o 2DGS-SN 8.5 8.6 47.5 54.1 50.3 33.52
w/o Merge 6.6 5.7 91.8 92.7 91.7 140.72
w/o Filter 6.8 5.8 90.9 92.9 91.3 44.37
Table 3: Ablation on the proposed components using ABC-NEF. w/o 2DGS-SN denotes SketchSplat trained with edge images detected by DexiNed, while the others are trained with 2DGS-SN.
Key Components.
[…]
EdgeGS 7.4 7.2 75.7 86.9 80.3 140.76
Ours (EdgeGS) 6.8 5.8 90.8 92.9 91.3 44.28
EMAP 8.8 7.9 63.5 70.4 66.3 48.71
Ours (EMAP) 6.9 6.2 88.8 90.8 89.3 31.70
Table 4: Ablation on initialization methods using ABC-NEF. Ours (EdgeGS) and Ours (EMAP) denote our method initialized with EdgeGS and EMAP, respectively. 2DGS-SN is chosen as 2D edge detector.
Initialization Methods.
In our main experiments, we use EdgeGS for initialization. However, we demonstrate SketchSplat can also achieve strong results when initialized with alternative methods like EMAP. As shown in Tab. 4, our approach consistently outperforms the initialization methods by a large margin and achieves comparable results across different initializations.
6 Conclusion
We present SketchSplat, a method for parametric 3D edge reconstruction from calibrated multi-view images. Unlike previous approaches that rely on 3D edge point reconstruction and edge fitting, SketchSplat directly optimizes edges through multi-view sketch splatting, ensuring accurate alignment with input 2D images. A series of topological operations are proposed to adaptively refine the edge structure during training, which reduces redundancy while maintaining high accuracy and completeness. Additionally, an improved 2D edge detection method is adopted to achieve higher reconstruction quality. Experiments show that SketchSplat achieves the state-of-the-art accuracy and completeness on the CAD dataset while maintaining compact results and efficient training speed.
Acknowledgements
Supplementary Material
Summary.
In this supplementary material, we elaborate on the following topics:
  • Sec. 7 provides details of our proposed edge detector 2DGS-SN, along with visual comparisons against existing 2D edge detectors.
  • Sec. 8 presents an alternative depth estimator for 2D edge detection.
  • Sec. 9 evaluates the robustness of our method under noisy initialization.
  • Sec. 10 discusses the necessity of the sketch filtering operation.
  • Metric definitions used in the main paper are detailed in Sec. 11.
  • Sec. 12 presents additional qualitative results, including 20 scenes from the ABC-NEF dataset.
  • In Sec. 13, we report quantitative results on the DTU dataset.
  • Sec. 14 compares our method with a feed-forward method.
  • Finally, Sec. 15 discusses the limitations of our approach.
7 Details of 2D Edge Detector
We show the accuracy of 2DGS-SN by comparing with PiDiNet [45] and DexiNed [34]. From Fig. 5 (b), we can see 2DGS-SN produces accurate edges, where the object boundary locates at the center of the predicted edge pixels. Both PiDiNet and DexiNed (Fig. 5 (c, d)) detect edges with slight offsets, positioning them far from the object boundaries. According to Tab. 1 in the main paper, this slight offset introduces significant multi-view inconsistency and ambiguity, resulting in degraded performance. This also explains why EdgeGS [8] achieve much higher F10 and F20 scores than EMAP [20], but lower F5 scores.
Figure 5: Visual comparison of different edge detection methods. Our 2DGS-SN provides more accurate edge detections results that align well with the object boundaries.
8 Alternative 2D Edge Detector
The core insight of 2DGS-SN is that geometric cues (e.g., depth and normal) offer higher pixel-level accuracy and are thus more robust than pure learning-based edge detectors [34, 45]. Based on this observation, we demonstrate that other geometry estimation methods such as DepthPro [6] can be alternative components of our 2DGS-SN. We replace 2DGS depth maps with the ones from DepthPro and test its performance. DepthPro is significantly faster than 2DGS, as it avoids per-scene optimization. However, its accuracy is lower due to unstable edge detection caused by multi-view inconsistent depth scales (see Tab. 6). In contrast, 2DGS [15] provides depth maps with a multi-view consistent scale, therefore a single threshold $t_{d}$ in Eq. 3 is sufficient to detect consistent edges. Choosing 2DGS (higher accuracy) vs. DepthPro (faster) is therefore a trade-off between efficiency and quality.
[…]
EdgeGS 2DGS-SN 7.4 7.2 75.7 86.9 80.3
Ours 2DGS-SN 6.8 5.8 90.8 92.9 91.3
EdgeGS DepthPro-SN 10.6 7.0 75.5 81.8 77.9
Ours DepthPro-SN 9.2 5.4 90.3 87.5 87.8
Table 6: Ablation on 2D edge detector. We ablate the depth prediction methods (2DGS [15], DepthPro [6]). 2DGS shows better performance since it produces multi-view consistent depth maps. All experiments use EdgeGS as initialization method.
9 Robustness to Initialization
To illustrate the robustness of our method, we inject Gaussian noise $\delta\sim\mathcal{N}(0,\sigma^{2})$ to the parameters of the edges initialized from EdgeGS. Fig. 6 and Tab. 7 show that our method performs robust even with noisy initializations and only degrades when initial edge lengths become excessively large relative to the scene size, which rarely occurs in practice. Exploring alternative initialization methods would be an interesting topic for future work.
[…]
EdgeGS - 7.4 7.2 75.7 86.9 80.3
Ours - 6.8 5.8 90.8 92.9 91.3
Ours 0.01 6.7 5.7 91.0 92.9 91.4
Ours 0.02 6.6 6.2 91.0 92.4 91.2
Ours 0.05 7.4 8.4 87.3 84.3 85.2
Ours 0.10 21.4 17.5 65.8 54.1 58.3
Table 7: Performance under Noisy Initialization. We present results with noisy initialization by adding Gaussian noise to the input edge parameters (on ABC-NEF dataset). Compared to the baseline EdgeGS, our method performs significantly better within a reasonable noise level ($\sigma<0.05$ for a $1m^{3}$ scene). All results use 2DGS-SN as edge detector.
Figure 6: Performance under noisy initialization. We add Gaussian noise $\delta\sim\mathcal{N}(0,\sigma^{2})$ to our initialization edges (from EdgeGS). On ABC-NEF, our method remains robust within reasonable noise levels but degrades when initial edge lengths become excessively large relative to the scene size $1m^{3}(\sigma>0.02)$, which is rare in our experiments.
10 Necessity of Sketch Filtering
In the main paper, the proposed filtering operations seem to have little effect on the metrics (Tab. 3), as the edges initialized from EdgeGS are already clean. However, under poor initialization, filtering becomes essential for improving accuracy. We demonstrate this using a pseudo noisy initialization scenario, where we add Gaussian noise to the initialized edge parameters with $\sigma=0.02$ and examine the effect of filtering. Tab. 8 shows that without filtering, the method suffers a noticeable accuracy drop and slightly higher completeness, confirming our assumption.
[…]
Ours - all 6.8 5.8 90.8 92.9 91.3
Ours 0.02 all 6.6 6.2 91.0 92.4 91.2
Ours 0.02 no filter 7.0 5.9 91.0 90.7 90.4
Table 8: Effectiveness of Filtering Operations under Noisy Initialization. We evaluate performance under noisy initialization by adding Gaussian noise to the input edge parameters on the ABC-NEF dataset. Without filtering, the method shows a clear drop in accuracy.
@@ 10 Necessity of Sketch Filtering
[…]
@@ 12 More Visual Results
13 Quantitative Evaluation on DTU dataset
Tab. 9 shows the evaluation results on six scenes in the DTU dataset [1]. We reproduce EdgeGS [8] using its released code and configuration (denoted as EdgeGS*) to initialize sketches for our method. As shown in Tab. 9, SketchSplat without topological operations achieves comparable results to EMAP [20] and slightly outperforms EdgeGS, demonstrating its effectiveness in correcting misalignment between 3D edges and 2D edge images. However, incorporating topological operations slightly reduces recall, as merging reduces redundant edges, whereas the pseudo ground-truth favors duplicated edges in P5 and R5 due to the thickness of the GT annotations (see Fig. 4 in the main paper).
14 Comparison to Feed-forward Baseline
We also compare against a feed-forward 3D line detection method. To evaluate methods like NerVE [58], we first reconstruct a point cloud from the input images (e.g., using 2DGS), then input it to the pretrained NerVE model to predict 3D edges. As shown in Tab. 10, NerVE performs significantly worse than our method and the baselines, suggesting that optimization-based methods remain more robust and generalize better, while feed-forward approaches are more sensitive to noisy inputs.
[…]
EMAP 2DGS-SN 8.8 7.9 63.5 70.4 66.3
EdgeGS 2DGS-SN 7.4 7.2 75.7 86.9 80.3
Ours 2DGS-SN 6.8 5.8 90.8 92.9 91.3
NerVE (PWL) 2DGS 18.0 35.7 40.5 62.3 48.4
NerVE (CAD) 2DGS 15.3 117.3 24.0 57.7 32.3
Table 10: Comparison with feed-forward method NerVE on ABC-NEF dataset. For NerVE, we train a 2DGS and extract surface points as input. It turns out that NerVE is sensitive to noise in the input and produces incomplete and inaccurate edges. Our method use EdgeGS as initialization.